22madison:

A large community college has professors and lecturers. The total number of faculty members is 136. The school reported that they have 6 professors for every 11 lecturers. How many professors does the community college employ?

snowflake0531:

What is the professor to lecturer ratio?

snowflake0531:

it says `6 professors for every 11 lecturers` So what is the ratio between the two

snowflake0531:

you can reply me if you do, if you don't, you can tell me:)

snowflake0531:

Welp, if you're not going to reply, I'm just going to go ahead lol the ratio is 6:11, because 6 professors=11 lecturers This means that if there were 17 in all, there are 6 professors

snowflake0531:

So, we have 6/17 then we have total number of faculty members is 136 so, we have to find, if there are 136 ppl, how many of those are professors we do 6/17 = x/136

snowflake0531:

\[\frac{ 6}{ 17} = \frac{ x }{ 136 }\]

snowflake0531:

We then cross multiply 6 * 136 = 17x 816=17x x=48

snowflake0531:

So, there are 48 professors

22madison:

sorry you typed so fast but anyway thank you

snowflake0531:

yw~